1. Acquired cash from the issue of common stock. 2. Paid cash to reduce the principal on a bank note 3. Sold land for cash at an amount equal to its cost. 4. Provided services to clients for cash 5. Paid utilities expenses with cash 6. Paid a cash dividend to the stockholders Required Indicate which of the above events occur in the equation by selecting the letter for increase, the letter D for decrease, /D for increase and decrease in a particular component, and NA for does not affect under each of the components of the accounting equation. The first event is shown as an example OLIVE ENTERPRISES Accounting Equation Stockholders Equity Event Number Assets Liabilities Common Stock Earnin + NA 4 6
Answer #1

Answer

OLIVE ENTERPRISES
Accounting equation
Stockholder's equity
Event number Assets = Liabilities + Common Stock + Retained earnings
1. I = NA + I + NA
2. D = D + NA + NA
3. NA = NA + NA + NA
4. I = NA + NA + I
5. D = NA + NA + D
6. D = NA + NA + D